CCDs -- Known issues
|Occasionally, I will update some
of the known problems (and fixes) for the MIKE ccd system.
Radioactive dewar window for MIKE_RED. The original window for the red ccd was a BK-7 glass that was mildly radioactive, producing lots of "cosmic ray hits" on the detector. The window was swapped for a fused silica version.
Dark current noise contribution. Both MIKE ccds are often used binned 2×2 for hour-long exposures. A dark current of just 1.0 e-/pix/hr (or 4 e-, binned 2×2) adds a noise term of 2 e-. This is greater than the 1.8 e- read noise of MIKE_BLUE, and would degrade the 3.5 e- read noise of MIKE_RED. Operating the red ccd at -125 °C and the blue ccd at -135 °C minimizes the dark current. Update: upgrading both the detectors to ccd44-82 devices has reduced the dark current to < 1e-/pix/hr.
Dropped frames. On both red and blue sides, the optical link of the ccd controller would occasionally lose sync, with the result that frames were getting dropped pretty frequently. That is, the ccd readout would sometimes hang after the exposure was complete, and the readout would abort. New dsp and pci cards with a more robust link have been designed and installed to fix the problem. Update: the new link has been more robust, and link errors are no longer occurring.
Unstable QE for CCID-20. The blue ccd has shown some signs of QE instability. The QE is apparently restored by oxygen flooding at room temperature, followed by pumping and cooling. This needs to be done every time the ccd is warmed up.
Low full well for CCID-20. For quite a while, the blue ccd full well was limited to about 8000 DN in 1×1 mode, and 16000 DN in 2×2 binning. Not so good. We have now figured out the combination of clock and bias voltages to restore the full well to 65K DN.
Update: the unstable QE prompted us to replace the CCID-20 with an E2V ccd44-82, which has a large full well (>100 ke-) and stable QE over the whole device.
Stuck bits on ADC. There have been a couple of cases where the ADC has shown stuck bits. Powering off and letting it cool down, then restarting doesn't seem to fix the problem. The ADC gets replaced.
CCD cleaning. As can be seen in this series of photos, there was some contamination on the red ccd which we attempted to clean off.
Greg Burley (firstname.lastname@example.org)
Ian Thompson (email@example.com)